# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v1

@@ 4.2 Closed-Loop RL Fine-Tuning
For the identified CBVs, we replace rule-based control with the IL pre-trained trajectory generation model (Section 3.2), which generates $N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories per CBV. The highest-scoring trajectory is selected for closed-loop execution. However, due to the covariate shift between open-loop training and closed-loop deployment, this approach often accumulates errors over time, resulting in unrealistic long-term behavior. To address this issue, prior work has explored RL [58, 59] and hybrid IL-RL approaches [21, 22] for closed-loop fine-tuning. However, these methods typically compute the loss only on the executed action or the best trajectory, overlooking the remaining candidates and neglecting the inherently multimodal nature of real-world driving behaviors. Motivated by the success of GRPO [20] in large language models fine-tuning, we propose to treat all the candidate trajectories $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N{\mathrm{ref}}\times N_{\mathrm{lon}}}$ as a group output. We then compute the corresponding trajectory-level group-relative reward =$\mathcal{R}=\{r_{i}\}_{i=1}^{N{\mathrm{ref}}\times N_{\mathrm{lon}}}$ and the group-relative advantage $\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}$ as follows:
$r_{i}=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\tau_{i}^{t})\right],\ \hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{std}(\mathcal{R})},\quad\forall i=1,\dots,G$ (6)
where $G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}$, and $\mathrm{StateWiseRM}$ is the user-defined state-wise reward model. With the group relative advantage $\mathcal{A}$, we derive the GRPO objective in our framework as:
@@
where ${\pi}_{\theta}$ is the current trajectory generation model, ${\pi}_{\mathrm{ref}}$ is the reference policy model, $\epsilon$ and $\beta$ are hyper-parameters, and $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ is the trajectory-level importance sampling ratio.
By leveraging IL pre-training to capture trajectory-level realism and multimodality, our closed-loop RL fine-tuning serves as an alignment mechanism that bridges the gap between user preferences and real-world data distributions, while preserving trajectory-level realism and multimodality. Building on this foundation, we remove the KL regularization term used in the original GRPO formulation, as the IL pre-trained distribution may significantly diverge from the user-preferred output. In such cases, enforcing constraints via KL divergence is unnecessary and may hinder alignment.
Removing the KL regularization introduces instability during fine-tuning. PPO-Clip [58] addresses this by constraining updates through a clipped importance sampling ratio, effectively limiting policy shifts to a trust region. However, this mechanism is inadequate in our group-relative setting. The original clipping mechanism is designed for on-policy learning, where only the executed action is corrected. In contrast, we evaluate an entire set of trajectory candidates, where the trajectory-level importance ratio $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ can become arbitrarily large, especially when $\pi_{\theta}(\tau_{i}|s)$ dominates $\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)$ and $\hat{A}_{i}<0$, the resulting product $\rho_{i}(\theta)\hat{A}_{i}$ yields large negative values, leading to high variance and unstable updates. To stabilize fine-tuning while preserving alignment with user-defined preferences, we adopt the dual-clip mechanism from Dual-Clip PPO [60, 61], where a constant $c>1$ acts as a lower bound on clipped negative advantages. The resulting RIFT objective is defined below (see Appendix A for theoretical analysis).
@@ 5 Experiment
This section systematically investigates the following research questions: Q1: How do the realism and controllability of interactive scenarios generated by RIFT compare to the baselines? Q2: How can the generated scenarios be leveraged for downstream tasks in autonomous driving? Q3: What are the contributions of each component within RIFT to the overall performance, and how well is controllability preserved under different user-specified reward settings? We first present the experimental setup, including the environment, baseline methods, and evaluation metrics. We then analyze the results to address these research questions.
@@ 5.1 Experiment Setups
Under the dual-stage AV-centered simulation framework, we adopt Pluto [53] as our trajectory generation model, owing to its strong performance and publicly available implementation. To ensure fairness during pre-training, we directly utilize the IL pre-trained checkpoint released by Pluto, which was pre-trained on the nuPlan dataset [11]. This setup enables a focused investigation of closed-loop fine-tuning. We conduct simulations in CARLA [62], building upon the Bench2Drive [46] framework to support comprehensive AV-centered closed-loop simulation and evaluation. Additional details regarding implementation, framework design, training protocols, and evaluation settings are provided in Appendix B.
@@
Baseline. To systematically evaluate the effectiveness of RIFT in closed-loop fine-tuning, we compare it against the following baselines (details are provided in Section B.5).
  • Pure RL/IL: Methods trained purely with either RL or IL without fine-tuning, including Pluto [53], FREA [14], FPPO-RS [14], and PPO [14].
  • RLFT/SFT: Methods that apply RL or supervised signals to fine-tune the pre-trained Pluto models, such as PPO-Pluto [58], REINFORCE-Pluto [59], GRPO-Pluto [20], and SFT-Pluto.
  • Hybrid: Methods that integrate both RL and supervised signals during fine-tuning, including RTR-Pluto [21] and RS-Pluto [22].
@@
All methods are fine-tuned on the same module—the trajectory scoring head of Pluto—to ensure fair comparisons and decouple trajectory-level realism from behavior-level reliability, as confirmed by the ablation studies in Section 5.4. Following the realism standards of the Sim Agent Challenge in WOSAC [34], we adopt a normal reward setting for all RL-based baselines, with the detailed configuration provided in Section B.6. Results under an aggressive reward configuration are also reported in Section 5.4.
Metrics. We extend the WOSAC evaluation framework [34], which evaluates behavior controllability and realism across three key categories—motion, agent interactions, and map adherence—to define our own set of evaluation metrics, as inspired by [63]. More details are provided in Section B.7.
  • Controllability: We measure infractions using Scenario Collision Per Kilometer (CPK) and Off-Road Rate (ORR)[41, 42], evaluate driving progress via Route Progress (RP), and assess driving safety with safety-critical metrics such as 2D-TTC[64] and ACT [65].
  • Realism: Quantifying realism in traffic scenarios remains a challenging problem. Inspired by prior work that evaluates realism from a distributional perspective [21, 34, 63]. We use the Shapiro–Wilk test (SW) [66, 67] to assess the normality of speed and acceleration, and Wasserstein Distance (WD) [68] to measure deviation from the target speed in the CARLA [62].
@@
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite [69] AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
Method Type Controllability Realism
Infraction Progress Safety Speed Acceleration
ORR $\downarrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ SW $\uparrow$ WD $\downarrow$ SW $\uparrow$
Pluto [53] IL 0.24 $\pm$ 0.15 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01
PPO [14] RL 9.17 $\pm$ 2.39 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02
FREA [14] RL 9.01 $\pm$ 2.09 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01
FPPO-RS [14] RL 8.60 $\pm$ 0.25 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03
SFT-Pluto SFT 0.06 $\pm$ 0.07 6.33 $\pm$ 2.23 780.48 $\pm$ 41.05 2.20 $\pm$ 1.64 2.12 $\pm$ 1.51 0.88 $\pm$ 0.02 6.01 $\pm$ 0.19 0.87 $\pm$ 0.02
RS-Pluto [22] SFT+RLFT 1.05 $\pm$ 0.31 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01
RTR-Pluto [21] SFT+RLFT 0.08 $\pm$ 0.09 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03
PPO-Pluto RLFT 0.07 $\pm$ 0.13 6.89 $\pm$ 3.19 683.57 $\pm$ 38.12 2.66 $\pm$ 1.50 2.60 $\pm$ 1.43 0.95 $\pm$ 0.01 4.96 $\pm$ 0.31 0.90 $\pm$ 0.02
REINFORCE-Pluto RLFT 1.37 $\pm$ 1.13 6.98 $\pm$ 0.86 813.70 $\pm$ 24.76 2.39 $\pm$ 1.64 2.30 $\pm$ 1.55 0.92 $\pm$ 0.01 5.63 $\pm$ 0.19 0.90 $\pm$ 0.02
GRPO-Pluto RLFT 0.10 $\pm$ 0.08 7.24 $\pm$ 4.04 892.65 $\pm$ 65.27 2.65 $\pm$ 1.44 2.61 $\pm$ 1.48 0.94 $\pm$ 0.04 4.96 $\pm$ 0.89 0.96 $\pm$ 0.00
RIFT-Pluto (ours) RLFT 0.36 $\pm$ 0.20 6.83 $\pm$ 2.62 995.33 $\pm$ 84.62 2.74 $\pm$ 1.30 2.71 $\pm$ 1.32 0.97 $\pm$ 0.01 4.46 $\pm$ 0.43 0.93 $\pm$ 0.01
@@ 5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenarios across different CBV methods, with results summarized in Table 1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower infraction rates, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite [69], which prioritizes safety by avoiding risky maneuvers.
@@
This conservative tendency is further highlighted in Figure 3, where supervised policies exhibit significantly lower speed and acceleration profiles. In contrast, RIFT strikes a more favorable balance between safety and interactivity. It achieves superior safety performance, as reflected by higher 2D-TTC and ACT scores, while avoiding the overly cautious behaviors typical of supervised approaches. As shown in Figure 3, RIFT demonstrates higher average speed and acceleration, indicating more interactive behavior, while maintaining realistic motion profiles.
Figure 4: Qualitative examples illustrating the robustness of RIFT across diverse AV-centered traffic scenarios. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
@@
Figure 5: Temporal comparisons illustrating RIFT’s superior performance over other baselines under AV-centered closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
@@
Qualitative Results. To further demonstrate the effectiveness of RIFT, we present simulation results across a range of traffic scenarios under diverse conditions, including varying weather (e.g., rain and fog) and traffic scenarios (e.g., merging and crossing), as illustrated in Figure 4. These results demonstrate the robustness and generalizability of RIFT in challenging environments. In addition, we compare closed-loop simulations against representative baselines, as shown in Figure 5. Baseline methods often suffer from trajectory instability or poor-quality generation in closed-loop settings. In contrast, RIFT consistently generates smooth, high-quality trajectories with better temporal consistency. Additional qualitative results are provided in Section D.3.
5.3 Realistic Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
In addressing Q2, we investigate the integration of RIFT into downstream tasks, specifically in the closed-loop AV evaluation. Building on the closed-loop evaluation framework [46, 70] in CARLA, we replace its hand-crafted scenarios with traffic flows generated by various CBV methods while retaining the same global route settings. To ensure fair and consistent evaluation, we adopt the standard AV evaluation metrics and introduce a new metric, Blocked Rate (BR), to measure situations where CBVs obstruct the AV. We evaluate two publicly available AV algorithms: the rule-based PDM-Lite [69] and the learning-based PlanT [71]. The results, presented in Table 3, demonstrate that under the RIFT-generated scenarios, AVs achieve near-optimal Driving Score (DS), Route Completion (RC), and Infraction Score (IS), along with the lowest BR. These findings highlight RIFT’s effectiveness in minimizing blocking while maintaining realism and interactivity, underscoring its superiority in closed-loop AV evaluation. Further details are available in Appendix C.

### v2

@@ 4.2 Closed-Loop RL Fine-Tuning
While effective in capturing trajectory-level realism and route-level controllability, the open-loop IL pre-training approach suffers from covariate shift between training and deployment, resulting in error accumulation and unrealistic long-term behavior in closed-loop settings. Prior works have addressed this issue through RL PPO ; sutton1999reinforce or hybrid IL–RL fine-tuning zhang2023rtr ; peng2024improving , but typically optimize only the executed trajectory, ignoring other candidates and thereby risking mode collapse.
To preserve multimodality during closed-loop fine-tuning, we adopt GRPO-style group-relative optimization shao2024grpo , which explicitly evaluates all candidate modalities and assigns greater relative advantages to those that better align with user-preferred styles. This approach enhances style-level controllability while maintaining multimodality. Considering closed-loop dynamics, we evaluate simulated rollouts rather than raw candidate trajectories to mitigate plan–rollout deviation. Specifically, given $G=N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories $\mathcal{T}=\{\tau_{i}\}_{i=1}^{G}$ for a particular CBV, we perform short-horizon forward simulation dauner2023tuplan_garage to obtain rollouts $\widetilde{\mathcal{T}}=\{\widetilde{\tau}_{i}\}_{i=1}^{G}$, considered jointly as a group. Each rollout is evaluated by a user-defined state-wise reward model $\mathrm{StateWiseRM}$, from which we compute the corresponding discounted returns $\mathcal{R}=\{r_{i}\}_{i=1}^{G}$ and derive the group-relative advantages $\mathcal{A}=\{\hat{A}_{i}\}_{i=1}^{G}$ as follows:
$r_{i}=\sum_{t=0}^{T}\gamma^{t}\left[\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t})\right],\ \hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{std}(\mathcal{R})},\quad\forall i=1,\dots,G$ (6)
Here, $\hat{A}_{i}$ measures each candidate rollout’s performance relative to the group, promoting high-return rollouts while preserving multimodality. Given the group-relative advantage $\hat{A}_{i}$, we derive the standard GRPO objective in our framework as:
@@
where ${\pi}_{\theta}$ is the current planning model, ${\pi}_{\mathrm{ref}}$ is the reference model, $\epsilon$ and $\beta$ are hyper-parameters, and $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ is the trajectory-level importance sampling ratio.
Due to covariate shift in closed-loop deployment, the trajectory scoring head—lacking reference-line priors—is more susceptible to distribution mismatch than the trajectory generation head. As a result, the pre-trained scoring head may produce unstable or misaligned results. In contrast, the generation head, anchored by route-level priors, remains robust, consistently generating realistic and multimodal trajectories. To preserve this capability while enhancing style-level controllability and mitigating covariate shift, we freeze the generation head and fine-tune only the scoring head toward user-preferred styles. In this setting, constraining the scoring head with KL regularization to the IL pre-trained model would anchor learning to a biased reference, thereby hindering adaptation. We therefore remove the KL term, allowing the scoring head to adapt freely toward user preferences while retaining the trajectory-level realism and route-level controllability established during IL pre-training. Moreover, closed-loop RL fine-tuning enables the model to recover from non-expert states encountered during deployment, further alleviating covariate shift in trajectory scoring.
Removing KL regularization can introduce instability during fine-tuning. PPO-Clip PPO mitigates this by constraining policy updates via a clipped importance sampling ratio, effectively restricting shifts to a trust region. However, this mechanism is insufficient in our group-relative setting: the original clipping targets only the executed action in policy learning, whereas we evaluate an entire set of trajectory candidates. Here, the trajectory-level importance ratio $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}|s)/{\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)}$ can grow arbitrarily large, when $\pi_{\theta}(\tau_{i}|s)$ dominates $\pi_{\theta_{\mathrm{old}}}(\tau_{i}|s)$ and $\hat{A}_{i}<0$, the product $\rho_{i}(\theta)\hat{A}_{i}$ may yield large negative values, causing high variance and unstable updates. To address this while maintaining alignment with user-preferred styles, we adopt the dual-clip mechanism from Dual-Clip PPO ye2020dual_clip ; gao2021learning , where a constant $c>1$ bounds clipped negative advantages from below. Building on the theoretical analysis in Appendix˜A, we define the resulting RIFT objective as:
@@ 5 Experiment
This section systematically addresses the following research questions: Q1: How does RIFT compare with representative baselines in terms of the realism and controllability of the generated traffic scenarios? Q2: How can the generated traffic scenario be effectively utilized to support downstream autonomous driving tasks? Q3: What are the contributions of each component within RIFT to the overall performance, and how well is style-level controllability preserved under different user-specified driving styles? We first describe the experimental setup, including the simulation environment, baseline methods, and evaluation metrics, followed by a detailed analysis of the results in response to the proposed research questions.
@@ 5.1 Experiment Setups
Under the dual-stage AV-centric simulation framework, we adopt Pluto cheng2024pluto as our planning model for its well-established performance and open-source implementation. To ensure fair comparison, we use the official IL pre-trained checkpoint provided by Pluto, trained on the nuPlan dataset caesar2021nuplan . Simulations are conducted in CARLA carla , leveraging Bench2Drive jia2024bench2drive to support AV-centric closed-loop simulation and evaluation. Implementation details, training protocols, and evaluation settings are described in Appendix˜B.
@@
Baseline. To systematically evaluate the effectiveness of RIFT in closed-loop fine-tuning, we compare it against the following baselines, with implementation details provided in Section˜B.5.
@@
All methods are fine-tuned on the same module—the scoring head—to ensure fair comparisons, while isolating style-level controllability from trajectory-level realism and route-level controllability, as confirmed by the ablation studies in Section˜5.4. Following the realism standards of the Sim Agent Challenge in WOSAC montali2023wosac , we adopt a normal style reward for all RL-based baselines, with details in Section˜B.7. Results under an aggressive style reward are also reported in Section˜5.4.
Metrics. Building on the WOSAC evaluation framework montali2023wosac , we categorize our evaluation metrics into three groups: kinematic metrics, interaction metrics, and map metrics. Kinematic metrics focus on distributional motion properties (S-SW, S-WD, A-SW), as inspired by chen2024review ; huang2025IntSim , with the absence of ground-truth trajectories in CARLA precluding displacement-based measures (e.g., ADE, FDE). Interaction metrics evaluate agent interactions through collision frequency (Collision Per Kilometer, CPK), driving efficiency (Route Progress, RP), and safety-critical measures, including 2D-TTC guo20232dttc and ACT venthuruthiyil2022act . Map metrics evaluate adherence to road geometry through the Off-Road Rate (ORR). Collectively, these metrics provide a comprehensive evaluation of realism and controllability in closed-loop simulation; detailed definitions are provided in Section˜B.8.
@@
Table 1: Comparison in Controllability and Realism. Metrics are evaluated under the PDM-Lite Beibwenger2024pdmLite AV setting across three random seeds, with the best and the second-best results highlighted accordingly.
Method Type Kinematic Metrics Interaction Metrics Map Metrics
S-SW $\uparrow$ S-WD $\downarrow$ A-SW $\uparrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ ORR $\downarrow$
Pluto cheng2024pluto IL 0.88 $\pm$ 0.01 5.81 $\pm$ 0.06 0.90 $\pm$ 0.01 5.06 $\pm$ 2.69 564.14 $\pm$ 114.41 2.50 $\pm$ 1.48 2.44 $\pm$ 1.39 0.24 $\pm$ 0.15
PPO chen2024frea RL 0.95 $\pm$ 0.01 4.45 $\pm$ 0.15 0.89 $\pm$ 0.02 13.95 $\pm$ 2.34 409.51 $\pm$ 30.38 2.59 $\pm$ 1.60 2.52 $\pm$ 1.57 9.17 $\pm$ 2.39
FREA chen2024frea RL 0.93 $\pm$ 0.01 5.10 $\pm$ 0.14 0.93 $\pm$ 0.01 30.42 $\pm$ 5.28 292.81 $\pm$ 68.54 2.71 $\pm$ 1.40 2.67 $\pm$ 1.41 9.01 $\pm$ 2.09
FPPO-RS chen2024frea RL 0.87 $\pm$ 0.01 5.80 $\pm$ 0.11 0.80 $\pm$ 0.03 21.39 $\pm$ 3.23 356.79 $\pm$ 26.19 2.55 $\pm$ 1.69 2.53 $\pm$ 1.68 8.60 $\pm$ 0.25
SFT-Pluto SFT 0.88 $\pm$ 0.02 6.01 $\pm$ 0.19 0.87 $\pm$ 0.02 6.33 $\pm$ 2.23 780.48 $\pm$ 41.05 2.20 $\pm$ 1.64 2.12 $\pm$ 1.51 0.06 $\pm$ 0.07
RS-Pluto peng2024improving SFT+RLFT 0.93 $\pm$ 0.00 5.40 $\pm$ 0.15 0.92 $\pm$ 0.01 4.11 $\pm$ 3.90 819.40 $\pm$ 74.07 2.27 $\pm$ 1.45 2.23 $\pm$ 1.43 1.05 $\pm$ 0.31
RTR-Pluto zhang2023rtr SFT+RLFT 0.85 $\pm$ 0.00 6.24 $\pm$ 0.16 0.81 $\pm$ 0.03 6.98 $\pm$ 2.59 481.60 $\pm$ 70.19 2.55 $\pm$ 1.60 2.47 $\pm$ 1.51 0.08 $\pm$ 0.09
PPO-Pluto RLFT 0.95 $\pm$ 0.01 4.96 $\pm$ 0.31 0.90 $\pm$ 0.02 6.89 $\pm$ 3.19 683.57 $\pm$ 38.12 2.66 $\pm$ 1.50 2.60 $\pm$ 1.43 0.07 $\pm$ 0.13
REINFORCE-Pluto RLFT 0.92 $\pm$ 0.01 5.63 $\pm$ 0.19 0.90 $\pm$ 0.02 6.98 $\pm$ 0.86 813.70 $\pm$ 24.76 2.39 $\pm$ 1.64 2.30 $\pm$ 1.55 1.37 $\pm$ 1.13
GRPO-Pluto RLFT 0.94 $\pm$ 0.04 4.96 $\pm$ 0.89 0.96 $\pm$ 0.00 7.24 $\pm$ 4.04 892.65 $\pm$ 65.27 2.65 $\pm$ 1.44 2.61 $\pm$ 1.48 0.10 $\pm$ 0.08
RIFT-Pluto (ours) RLFT 0.97 $\pm$ 0.01 4.46 $\pm$ 0.43 0.93 $\pm$ 0.01 6.83 $\pm$ 2.62 995.33 $\pm$ 84.62 2.74 $\pm$ 1.30 2.71 $\pm$ 1.32 0.36 $\pm$ 0.20
@@ 5.2 Realistic and Controllable Traffic Scenario Generation (Q1)
Main Results. To address Q1, we evaluate the controllability and realism of the generated scenario across different CBV methods, with results summarized in Table˜1. RIFT consistently outperforms all baselines in both aspects across most settings. While supervised learning methods achieve slightly lower CPK and ORR, this improvement is primarily due to their inherently conservative behavior, derived from the expert PDM-Lite Beibwenger2024pdmLite , which prioritizes safety by avoiding risky maneuvers.
@@
This conservative tendency is further highlighted in Figure˜3, where supervised policies exhibit significantly lower speed and acceleration profiles. In contrast, RIFT strikes a more favorable balance between safety and interactivity. It achieves superior safety performance, as reflected by higher 2D-TTC and ACT scores, while avoiding the overly cautious behaviors typical of supervised approaches. As shown in Figure˜3, RIFT demonstrates higher average speed and acceleration, indicating more interactive behavior, while maintaining realistic motion profiles.
@@
Figure 4: Temporal comparisons illustrating RIFT’s superior performance over other baselines under AV-centric closed-loop simulation. CBV is marked in purple, AV (PDM-Lite) is in red, and BVs are in blue.
@@
Qualitative Results. To further demonstrate the effectiveness of RIFT, we compare closed-loop simulations against representative baselines, as shown in Figure˜4. Baseline methods often suffer from unstable or low-quality trajectory selection in closed-loop settings, whereas RIFT consistently selects smooth, high-quality trajectories with superior temporal consistency. Further qualitative examples are presented in Section˜D.3.
5.3 Generated Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
To address Q2, we assess the suitability of traffic scenarios generated by different CBV methods for closed-loop AV evaluation. Following KING king , we adopt PDM-Lite Beibwenger2024pdmLite —a rule-based planner with privileged access—as a reference to evaluate two key scenario properties: feasibility, measured by Driving Score (DS), and naturalness, captured by our proposed Blocked Rate (BR). A high DS indicates that the AV can reliably complete the scenario, while a low BR reflects realistic interactions without excessive obstruction from surrounding vehicles. Together, DS and BR offer a principled basis for evaluating scenario quality.
To further assess each scenario’s ability to reveal weaknesses in learning-based planners, we compare the performance of PlanT Renz2022PlanT , UniAD hu2023_uniad , and VAD jiang2023vad against PDM-Lite. As these models are more sensitive to subtle or adversarial interactions, informative scenarios should induce noticeable performance drops. As shown in Table˜3, traffic generated by RIFT achieves the highest DS and lowest BR under PDM-Lite, while also causing the most significant degradation in all learning-based planners. These results demonstrate that RIFT produces realistic and well-structured scenarios that are effective at exposing the limitations of modern AV systems. See Appendix˜C for further details.